# TAMER: Tree-Aware Transformer for Handwritten Mathematical Expression Recognition

### v1

@@ Introduction
Table 2 presents TAMER’s performance on the HME100K dataset. Our approach outperforms CoMER and NAMER across all metrics, but slightly lags behind ICAL. Nonetheless, TAMER significantly surpasses ICAL in the syntactical accuracy of the predicted expressions. Further details can be found in the next section.
@@ Appendix B Inference Speed
As shown in Table  4, we compared the inference speed of TAMER with the baseline CoMER on a single NVIDIA 2080Ti GPU. To standardize the comparison, we used images from the CROHME dataset with sizes close to 16x256 and resized them uniformly to 16x256. For calculating frames per second (FPS), we set the batch size to 4 and conducted 25 iterations, totaling 100 images for inference, and averaged the results to determine the FPS. To ensure authenticity during inference, the model input consisted solely of the expression image and its corresponding image mask. It can be seen that the introduction of the Tree-Aware Module (TAM) results in a slight increase in the model’s parameter count and, due to the inclusion of tree structure prediction scoring mechanism, a slight decrease in FPS.
Method Params (M) FLOPs (G) FPS
CoMER 6.39 18.81 0.575
TAMER 8.23 19.79 0.384

### v2

@@ Introduction
Table 2 showcases the performance of TAMER on the HME100K dataset, where it exceeds the capabilities of both CoMER and NAMER across all assessed metrics. Compared to methods that typically incorporate auxiliary tasks, ICAL includes an additional Fusion Module, which effectively leverages auxiliary task information during training. Note that TAMER can also be easily extended with this kind of extra module. Consequently, we have included results for TAMER equipped with the Fusion Module. As indicated in the table, TAMER with the Fusion Module achieves SOTA performance on the HME100K dataset.
@@ Conclusion
Acknowledgements
This work is supported by the projects of National Natural Science Foundation of China (No. 62376012) and Beijing Science and Technology Program (Z231100007423011), which is also a research achievement of State Key Laboratory of Multimedia Information Processing and Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology).
@@ Appendix B Inference Speed
As shown in Table  4, we compared the inference speed of TAMER with the baseline CoMER on a single NVIDIA 2080Ti GPU. To standardize the comparison, we used images from the CROHME dataset with sizes close to 64x256 and resized them uniformly to 64x256. For calculating frames per second (FPS), we set the batch size to 1 and conducted 100 iterations, totaling 100 images for inference, and averaged the results to determine the FPS. To ensure authenticity during inference, the model input consisted solely of the expression image and its corresponding image mask. Our findings show that introducing the Tree-Aware Module (TAM) in TAMER results in a moderate increase in parameter count (from 6.39 M to 8.23 M) and a slight rise in computational complexity (GFLOPs increase from 18.81 to 19.79). The frames per second (FPS) decreased from 11.13 to 6.75 mainly due to the inclusion of the Tree Structure Prediction Scoring Mechanism. By omitting this tree scoring mechanism, TAMER achieves an FPS of 9.45.
Method Params (M) GFLOPs FPS ExpRate
CoMER 6.39 18.81 11.13 58.38
TAMER 8.23 19.79 6.75 61.23
w/o tree scoring 8.23 19.79 9.45 60.39